# Self-supervised pretraining of vision transformers for animal behavioral analysis and neural encoding

### v1

@@ B.1 Architecture
We selected the “base” ViT-B/16 architecture over other ViT variants for two reasons: it is expressive enough to capture rich frame-level information, while remaining computationally efficient for training and inference on long videos.
Our architecture differs from the standard ViT in its use of a nonlinear projector for the contrastive loss. While ViC-MAE employs a pooled attention layer to transform patch embeddings into a 768-dimensional vector for their contrastive loss, we take a different approach. We use the standard CLS token as our global image representation rather than pooled patch embeddings. This CLS token passes through a nonlinear projector with four components: a linear layer, Batch Norm (necessary for stable training, Fig. 5), ReLU activation, and a final linear layer.
@@ B.2 Training
[…]
Frame selection strategy
Animal behavior videos often contain extended periods of inactivity or repetitive behaviors. Pretraining ViT models on all available frames would capture redundant information and increase computation time unnecessarily. Instead, we focus on extracting diverse frames that showcase distinct poses (to optimize the masked autoencoding loss) while preserving meaningful temporal relationships in local neighborhoods (to leverage the temporal contrastive loss). Our approach begins by downsampling all video frames to 32 $\times$ 32 pixels and calculating motion energy, defined as the absolute pixel-wise differences between consecutive frames. We eliminate frames in the bottom 50th percentile of motion energy, retaining only those with significant movement. We then apply k-means clustering to the remaining downsampled frames, with the number of clusters matching our target number of anchor frames per video (e.g., 600). For each cluster, we select the frame closest to the cluster center, along with its immediate predecessor and successor in time, which serve as positive examples for the contrastive loss (for a total of, e.g., 1800 frames per video). This methodology ensures a high-quality, diverse dataset for efficient pretraining. We find this method outperforms a random frame selection strategy in the neural encoding task (Table 4).
Table 4: Frame selection strategy ablation. We pretrain models using either a random frame selection strategy (“Random”) or the PCA+k-means strategy described above (“Selected”). We evaluate the representations of these models using zero-shot performance on the neural encoding task using the bits per spike (BPS) metric. We report the mean and standard deviation of BPS across five test sessions.
@@
Frame sampling strategy during training
Once we have a diverse set of training frames, we must construct batches during training. As stated in Sec. 3, for a batch of size $B$ we randomly select $B/2$ anchor frames, which can originate from any and all videos. For Beast, each anchor frame $\mathbf{x}_{t}^{v}$ is paired with a positive frame randomly selected from $\mathbf{x}_{t\pm 1}^{v}$. All other frames serve as negative frames, including frames from the same video. Due to the frame selection strategy described above, even frames from the same video will be visually distinct and not interfere with the contrastive loss. This batch construction procedure is distinct from ViC-MAE, which allows any two frames from the same video to be a positive pair, while only frames from different videos are negative pairs. We find our approach outperforms the ViC-MAE approach in the neural encoding task (Table 5). This sampling strategy only applies to Beast; the ViT-M models do not contain the contrastive loss, and we only train them with the anchor frames.
@@
Data augmentation
The default data augmentation procedure [15] applies a random resized crop to $244\times 244$ pixels with a crop ratio between 0.2 and 1.0, followed by a random horizontal flip with probability 50%. We also explore an extended augmentation strategy that adds further random transformations: rotation up to 45 degrees and color jittering (brightness=0.4, contrast=0.4, saturation=0.4, hue=0.1), in addition to the crop and flip. We compare the performance of Beast using both the default and the extended augmentation strategy. The additional augmentations achieve performance similar to the default setting (Table 6), so we use the default augmentation throughout the paper.
@@ B.3 Hyperparameter details
The Beast objective combines two losses: the reconstruction loss and the contrastive loss. During the early stages of training, it is important for the model to focus on accurately reconstructing the input, so the reconstruction loss should dominate. As training progresses and the model learns to capture low-level pixel structure, the contrastive loss gradually becomes more important. It acts as a regularizer, encouraging the model to learn higher-level temporal representations rather than overfitting to local pixel patterns.
The weighting factor for the contrastive loss, $\lambda$, plays a crucial role during pretraining; too large and the model will not reconstruct the input well; too small and the model does not reap its regularizing benefits. Through hyperparameter tuning based on neural encoding performance (on the validation set), we set $\lambda=0.03$ for all datasets, except CRIM13, where we set $\lambda=0.01$. This choice ensures the reconstruction loss is emphasized in the early phases of training, while the contrastive loss naturally takes over as reconstruction error decreases toward the end, eliminating the need for an annealing schedule.
@@ C.1 Feature representations
Whisker pad motion energy
We localize the whisker pad in the IBL dataset using anchor keypoints on the nose and eye. We then compute the motion energy of the whisker pad as the absolute pixel-wise differences between consecutive frames, resulting in a one-dimensional representation at each time point [30].
Principal Component Analysis
@@
CEBRA
CEBRA [55] is a contrastive learning approach that provides a baseline which is complementary to the ViT-M models pretrained on ImageNet with a masked autoencoding objective. Similar to our PCA approach, we train an individual unsupervised CEBRA model for each session using the convolutional neural network option and the default offset10-model.
ViT-M variants
@@
Beast variants
@@ C.2 Reduced rank regression
IBL
For the IBL dataset, we followed the Reduced Rank Regression (RRR) setup described in [51]. We trained all models using the L-BFGS optimizer and set the rank constraint to 3. To denoise the neural signals, we applied a 1-dimensional smoothing filter to the neural activity. The hyperparameter search (Sec. C.4) was conducted over the ranges specified in Table 7.
@@
Facemap
For the Facemap dataset, we followed the setup described in [63], using the implementation provided in the official Facemap repository. We adopted the default rank of 32 as used in the original implementation. The Lambda parameter refers to the regularization strength, which we set to a relatively low value to avoid over-penalizing the weights. The output dimensionality was set to 128, corresponding to the number of neural principal components used in the model. Model parameters are estimated via a closed-form least squares approach. The hyperparameters we used are specified in Table 8.
@@ C.3 Temporal convolution network
We used the same implementation of the Temporal Convolution Network (TCN) to process frame embeddings for both the IBL and Facemap datasets, based on the official Facemap repository. The convolutional kernel operates along the temporal dimension of the input data. The TCN model was trained for 300 epochs using the AdamW optimizer, with learning rate decimation (multiplied by 0.1) applied at epochs 120 and 200. The hyperparameter search (Sec. C.4) was conducted over the ranges specified in Table 9.
@@ C.4 Hyperparameter selection
For the IBL data, we divided the trials into train (80%), validation (10%), and test (10%) sets. For the Facemap data we followed the experimental setup as described in the original study [63]: the session is split into ten blocks; the first 75% of each block is assigned to the training set; the following 3 seconds are excluded to remove data leakage due to autocorrelation in behavior and neural activity; and the final set of frames from the block are assigned to the test set. There is no validation set. For both the RRR and TCN models, we conducted hyperparameter searches separately for each feature type to identify the best-performing configurations. Specifically, we performed 30 runs of randomly selected hyperparameters per model type, evaluating performance on the validation set (IBL) or test set (Facemap) using an evaluation metric specific to each dataset: bits per spike (BPS) for IBL and variance explained for Facemap. The only exception was the RRR model for Facemap, where the parameters were fixed according to the original implementation. We select the model with the best performance on the validation (IBL) or test (Facemap) set, and report results on the test set.
@@ C.5 Patch embeddings
The output of the ViT encoder consists of a CLS token embedding $z_{\text{CLS}}\in\mathbb{R}^{D}$ and patch token embeddings $z\in\mathbb{R}^{N\times D}$, where $D=768$ is the embedding dimension and $N$ is the total number of image patches ($N=196$ for a $224\times 224$ input frame). We compare the CLS token and patch embeddings as inputs for neural encoding (see implementation details in E.4), and find that the CLS token outperforms the patch embeddings (Table 10). Given its lower dimensionality and superior results, we adopt the CLS token representation for all subsequent encoding tasks.
Table 10: Comparison of CLS and patch embeddings with a TCN encoder.
@@ C.7 Extended neural encoding results
Table 11: Neural encoding results across feature types and models. All ViT-based models use a frozen backbone. “IN” refers to a model pretrained with ImageNet weights; “IN+PT” refers to models that are initialized with ImageNet-pretrained weights then further pretrained on experiment-specific data; “+FT” refers to models that are initialized with pretrained weights based on what comes before “+” then fine-tuned on individual sessions.
Features IBL IBL-whisker Facemap RRR TCN RRR TCN RRR TCN Keypoints $0.173\pm 0.029$ $0.271\pm 0.054$ $0.224\pm 0.047$ $0.403\pm 0.077$ Motion energy $0.087\pm 0.023$ $0.117\pm 0.028$ PCA $0.266\pm 0.054$ $0.312\pm 0.078$ $0.220\pm 0.038$ $0.281\pm 0.065$ $0.177\pm 0.064$ $0.407\pm 0.090$ CEBRA $0.245\pm 0.036$ $0.295\pm 0.049$ $0.209\pm 0.265$ $0.265\pm 0.034$ $0.143\pm 0.046$ $0.278\pm 0.064$ ViT-M (IN) $0.201\pm 0.070$ $0.325\pm 0.091$ $0.142\pm 0.051$ $0.307\pm 0.068$ $0.254\pm 0.061$ $0.446\pm 0.100s$ ViT-M (IN+PT) $0.182\pm 0.071$ $0.334\pm 0.098$ $0.156\pm 0.032$ $0.316\pm 0.073$ ViT-M (IN+FT) $0.242\pm 0.089$ $0.349\pm 0.106$ $0.219\pm 0.048$ $0.336\pm 0.075$ $0.260\pm 0.051$ $0.461\pm 0.099$ ViT-M (IN+PT+FT) $\mathbf{0.293\pm 0.082}$ $0.351\pm 0.106$ $0.244\pm 0.042$ $0.335\pm 0.092$ Beast (IN+PT) $0.277\pm 0.076$ $0.337\pm 0.103$ $0.138\pm 0.029$ $0.317\pm 0.083$ Beast (IN+FT) $0.276\pm 0.088$ $0.351\pm 0.106$ $\mathbf{0.251\pm 0.051}$ $\mathbf{0.337\pm 0.088}$ $\mathbf{0.281\pm 0.054}$ $\mathbf{0.464\pm 0.089}$ Beast (IN+PT+FT) $0.291\pm 0.087$ $\mathbf{0.352\pm 0.106}$ $0.243\pm 0.048$ $0.335\pm 0.079$

### v2

@@ B.1 Architecture
We selected the “base” ViT-B/16 architecture over other ViT variants because it is expressive enough to capture rich frame-level information while remaining computationally efficient for training and inference on long videos. While ViC-MAE employs a pooled attention layer to transform patch embeddings into a 768-dimensional vector that serves as the global image representation for the temporal contrastive loss, we use the standard CLS token (Table 7). This CLS token is passed through a nonlinear projector consisting of four components: a linear layer, Batch Norm (necessary for stable training, Fig. 5), ReLU activation, and a final linear layer.
@@
The Beast backbone is initialized using weights pretrained on ImageNet with a masked autoencoding loss. This model has exceptional zero-shot performance on neural encoding, outperforming all other baselines, and is further improved with domain-specific pretraining (Table 1). We find other pretrained backbones–DINOv2 [Oquab et al., 2023] and CLIP [Radford et al., 2021]–also have strong zero-shot performance, but do not significantly improve upon MAE+ImageNet (Table 13), indicating this is a reasonable pretrained backbone from which to start our own domain-specific pretraining (see Fig. 9 and Table 15 for similar results on pose estimation and action segmentation, respectively).
Beast incorporates time through its temporal contrastive loss, which efficiently captures information across frames. We explored the performance of VideoMAE [Tong et al., 2022], a related video model pretrained on Kinetics-400 [Kay et al., 2017]. We found that the frozen VideoMAE backbone does outperform the frame-based ViT-MAE on the neural encoding task (Table 13). Interestingly Beast, which applies additional domain-specific pretraining to ViT-MAE, still outperforms VideoMAE. This raises the intriguing question of whether further domain-specific pretraining of VideoMAE could surpass Beast performance. We view this as an important direction for future work.
@@ B.2 Training
We discuss frame selection and sampling strategies, data augmentations, and global pooling strategies below. We apply the MAE loss uniformly across all frame types (anchor, positive, negative), whereas ViC-MAE only applies the MAE loss to anchor frames. The global batch size is set to 2048, distributed across 8 Nvidia A40 GPUs. We use the AdamW optimizer with a weight decay of 0.05. The learning rate is scheduled using PyTorch’s OneCycleLR scheduler, with a base learning rate of $5\times 10^{-5}$. The maximum learning rate is computed as $\text{max\_lr}=\text{base\_lr}\times\frac{\text{global\_batch\_size}}{256}$, with pct_start set to 0.15 and div_factor set to 10. We train all models for 800 epochs.
Frame selection strategy.
Animal behavior videos often contain extended periods of inactivity or repetitive behaviors. Pretraining ViT models on all available frames would capture redundant information and increase computation time. Instead, we extract diverse frames exhibiting distinct poses (to optimize the masked autoencoding loss) while preserving local temporal relationships (to leverage the temporal contrastive loss). We first downsample all frames to 32 $\times$ 32 pixels and calculate motion energy, defined as the absolute pixel-wise differences between consecutive frames. We eliminate frames in the bottom 50th percentile of motion energy, retaining only those with significant movement. We then apply $k$-means clustering to the remaining downsampled frames, with the number of clusters matching our target number of anchor frames per video (e.g., 600). For each cluster, we select the frame closest to the cluster center, along with its immediate predecessor and successor in time, which serve as positive examples for the contrastive loss (for a total of, e.g., 1800 frames per video). This methodology ensures a high-quality, diverse dataset for efficient pretraining and outperforms random frame selection in the neural encoding task (Table 4).
Table 4: Frame selection strategy ablation. We pretrain models using either random selection (“Random”) or the PCA+$k$-means strategy described above (“Selected”). We evaluate the representations of these models using zero-shot performance on the neural encoding task using the bits per spike (BPS) metric. We report the mean and standard deviation of BPS across five test sessions.
@@
Frame sampling strategy during training.
Once we have a diverse set of training frames, we must construct batches during training. As stated in Sec. 3, for a batch of size $B$ we randomly select $B/2$ anchor frames, which can originate from any and all videos. For Beast, each anchor frame $\mathbf{x}_{t}^{v}$ is paired with a positive frame randomly selected from $\mathbf{x}_{t\pm 1}^{v}$. All other frames serve as negative frames, including frames from the same video. Due to the frame selection strategy described above, even frames from the same video will be visually distinct and not interfere with the contrastive loss (Fig. 6). This batch construction procedure is distinct from ViC-MAE, which allows any two frames from the same video to be a positive pair, while only frames from different videos are negative pairs. We find our approach outperforms the ViC-MAE approach in the neural encoding task (Table 5). This sampling strategy only applies to Beast; the ViT-M models do not contain the contrastive loss, and we only train them with the anchor frames.
@@
Figure 6: Temporal contrastive loss visualization. UMAP embeddings [McInnes et al., 2018] of the first 32 principal components (PCs) of all anchor frames used for pretraining on the IBL dataset. Each anchor frame is colored by the session it was sampled from. The left and right columns show example frames drawn from the same session and from different sessions, respectively. The sampled frames are largely visually distinct, even frames next to each other in UMAP space (e.g., frames B and C, or D and E, on the left-hand side).
Data augmentation.
The default data augmentation procedure [He et al., 2022] applies a random resized crop to $244\times 244$ pixels with a crop ratio between 0.2 and 1.0, followed by a random horizontal flip with probability 50%. We explored an extended augmentation strategy that adds further transformations: rotation up to 45 degrees and color jittering (brightness=0.4, contrast=0.4, saturation=0.4, hue=0.1), in addition to the crop and flip. We compare the performance of Beast using both the default and the extended augmentation strategy. The additional augmentations achieve performance similar to the default setting (Table 6), so we use the default throughout the paper.
@@
Pooling strategy.
The CLS token serves as a global frame representation, effectively pooling information across all spatial positions into a single latent vector. To explore alternative pooling strategies during pretraining, we conducted additional experiments by pretraining models using mean pooling and attention pooling of the patch embeddings, then evaluating performance on the neural encoding task (Table 7). The ablation results clearly demonstrate that the CLS token is the most effective aggregation method for pretraining.
Table 7: Pooling ablation. We pretrain models using either the CLS token, or mean or attention pooling of the patch embeddings, to aggregate information across the image for the temporal contrastive loss. We evaluate the representations of these models using zero-shot performance on the neural encoding task using the bits per spike (BPS) metric. We report the mean and standard deviation of BPS across five test sessions.
Method IBL IBL-whisker
TCN RRR TCN RRR
Beast (mean pooling) $0.321\pm 0.104$ $0.159\pm 0.080$ $0.302\pm 0.082$ $0.128\pm 0.025$
Beast (attention pooling) $0.323\pm 0.100$ $0.141\pm 0.076$ $0.307\pm 0.072$ $0.130\pm 0.031$
Beast (CLS token) $\mathbf{0.337\pm 0.103}$ $\mathbf{0.277\pm 0.076}$ $\mathbf{0.317\pm 0.083}$ $\mathbf{0.138\pm 0.029}$
@@ B.3 Hyperparameter details
The Beast objective combines reconstruction and contrastive losses. During the early stages of training, it is important for the model to focus on accurately reconstructing the input, so the reconstruction loss should dominate. As the model learns to capture low-level pixel structure, the contrastive loss gradually becomes more important. It acts as a regularizer, encouraging the model to learn higher-level temporal representations rather than overfitting to local pixel patterns.
The weighting factor for the contrastive loss, $\lambda$, plays a crucial role during pretraining; too large and the model will not reconstruct the input well; too small and the model does not reap its regularizing benefits. Through hyperparameter tuning based on neural encoding performance (on the validation set), we set $\lambda=0.01$ for all datasets, except IBL and mirror-mouse, where we set $\lambda=0.02$. This choice ensures the reconstruction loss is emphasized in the early phases of training, while the contrastive loss naturally takes over as reconstruction error decreases toward the end, eliminating the need for an annealing schedule.
@@
B.4 Computational efficiency comparison
To compare the computational efficiency of image (ViT-M, Beast) versus video models (VideoMAE), we recorded three metrics for each model: runtime (ms per batch), Giga Floating Point Operations per Second (GFLOPS), and memory required for a forward pass with batch size one. All experiments were run using the fvcore library on a single A100 GPU. Since runtime varies across batches while GFLOPS and memory are deterministic, we report mean and standard deviation of runtime across 32 batches. We benchmark two modes: “pretrain”, which uses patch masking (0.75 for image models, 0.9 for VideoMAE, following optimal settings from the respective papers); and “finetune”, which omits patch masking as relevant for our downstream tasks.
[…]
Table 8: Computational efficiency comparison. We benchmark runtime (mean and standard deviation across 32 batches), GFLOPS, and memory usage for Beast, ViT-M, and VideoMAE during pretraining (with patch masking) and finetuning (without patch masking). All measurements are for a forward pass with batch size of one on a single A100 GPU.
Model Runtime (ms/batch) GFLOPS Memory (MB)
Finetune Pretrain Finetune Pretrain Finetune Pretrain
ViT-M $5.43\pm 0.70$ $7.6\pm 2.43$ $17.59$ $9.80$ $369.0$ $332.72$
Beast $6.21\pm 0.94$ $7.9\pm 2.26$ $17.59$ $9.80$ $409.1$ $333.40$
VideoMAE $71.51\pm 3.50$ $20.16\pm 2.47$ $199.49$ $35.24$ $2955.1$ $831.61$
B.5 Feature visualizations
We computed PCA on patch embeddings from 100 randomly sampled frames in the test set and visualized the first three components as RGB channels of two example frames. Compared to DINOv2–whose embeddings emphasize broader semantic structures–the Beast pre-trained model captures finer-grained details that are critical for neural encoding, pose estimation and behavior segmentation. In contrast, the ViT-M model pretrained only on ImageNet produces patch embeddings that appear noisier and less structurally coherent.
Figure 7: Visualization of the first three PCA components across models.
@@ C.1 Feature representations
Whisker pad motion energy.
We localize the whisker pad in the IBL dataset using anchor keypoints on the nose and eye. We then compute the motion energy of the whisker pad as the absolute pixel-wise differences between consecutive frames, resulting in a one-dimensional representation at each time point [IBL et al., 2022].
Principal Component Analysis.
@@
CEBRA.
CEBRA [Schneider et al., 2023] is a contrastive learning approach that provides a baseline which is complementary to the ViT-M models pretrained on ImageNet with a masked autoencoding objective. Similar to our PCA approach, we train an individual unsupervised CEBRA model for each session using the convolutional neural network option and the default offset10-model.
DINOv2.
For the DINOv2 model [Oquab et al., 2023], we extract the CLS embedding for each frame using a frozen pretrained backbone, and use these as input to the encoding models. We did not pretrain this model ourselves, but rather used the model checkpoint available at https://huggingface.co/facebook/dinov2-base.
ViT-M variants.
@@
ViT-C variants.
For the ViT-C (IN+PT) model, we start with the ViT-M (IN) weights then continue pretraining on dataset-specific frames with the temporal contrastive loss only. Therefore there is no ViT-C (IN) variant. We did not perform additional session-specific fine-tuning, so there is no ViT-C (IN+PT+FT) variant.
Beast variants.
@@ C.2 Reduced rank regression
IBL.
For the IBL dataset, we followed the Reduced Rank Regression (RRR) setup described in [Posani et al., 2025]. We trained all models using the L-BFGS optimizer and set the rank constraint to 3. To denoise the neural signals, we applied a 1-dimensional smoothing filter to the neural activity. The hyperparameter search (Sec. C.4) was conducted over the ranges specified in Table 9.
@@
Facemap.
For the Facemap dataset, we followed the setup described in [Syeda et al., 2024], using the implementation provided in the official Facemap repository. To deal with different neural and behavioral sampling rates, this model first resamples the behavioral timestamps to match the neural timestamps, and then fits a Reduced Rank Regression model using low-rank SVD. We adopted the default rank of 32 as used in the original implementation. The Lambda parameter refers to the regularization strength, which we set to a relatively low value to avoid over-penalizing the weights. The output dimensionality was set to 128, corresponding to the number of neural principal components used in the model. Model parameters are estimated via a closed-form least squares approach. The hyperparameters we used are specified in Table 10.
@@ C.3 Temporal convolution network
We used the same implementation of the Temporal Convolution Network (TCN) to process frame embeddings for both the IBL and Facemap datasets, based on the official Facemap repository. The convolutional kernel operates along the temporal dimension of the input (behavioral) data. To deal with different neural and behavioral sampling rates, this model resamples the resulting latent representation at neural timestamps using nearest-neighbor indexing. The TCN model was trained for 300 epochs using the AdamW optimizer, with learning rate decimation (multiplied by 0.1) applied at epochs 120 and 200. The hyperparameter search (Sec. C.4) was conducted over the ranges specified in Table 11.
@@ C.4 Hyperparameter selection
For the IBL data, we divided the trials into train (80%), validation (10%), and test (10%) sets. For the Facemap data we followed the experimental setup as described in the original study [Syeda et al., 2024]: the session is split into ten blocks; the first 75% of each block is assigned to the training set; the following 3 seconds are excluded to remove data leakage due to autocorrelation in behavior and neural activity; and the final set of frames from the block are assigned to the test set. There is no validation set. For both the RRR and TCN models, we conducted hyperparameter searches separately for each feature type to identify the best-performing configurations. Specifically, we performed 30 runs of randomly selected hyperparameters per model type, evaluating performance on the validation set (IBL) or test set (Facemap) using an evaluation metric specific to each dataset: bits per spike (BPS) for IBL and variance explained for Facemap. The only exception was the RRR model for Facemap, where the parameters were fixed according to the original implementation. We select the model with the best performance on the validation (IBL) or test (Facemap) set, and report results on the test set.
@@ C.5 Patch embeddings
The output of the ViT encoder consists of a CLS token embedding $z_{\text{CLS}}\in\mathbb{R}^{D}$ and patch token embeddings $z\in\mathbb{R}^{N\times D}$, where $D=768$ is the embedding dimension and $N$ is the total number of image patches ($N=196$ for a $224\times 224$ input frame). We compare the CLS token and attention-pooled patch embeddings as inputs for neural encoding (see implementation details in E.4), and find that the CLS token outperforms the patch embeddings (Table 12). Given its lower dimensionality and superior results, we adopt the CLS token representation for all subsequent encoding tasks.
Table 12: Comparison of CLS and patch embeddings with a TCN encoder. We report the mean and standard deviation of BPS across five test sessions.
@@ C.7 Extended neural encoding results
Table 13: Neural encoding results across feature types and models. All ViT-based models use CLS tokens from a frozen backbone. “IN” refers to a model pretrained with ImageNet weights; “IN+PT” refers to models that are initialized with ImageNet-pretrained weights then further pretrained on experiment-specific data; “+FT” refers to models that are initialized with pretrained weights based on what comes before “+” then fine-tuned on individual sessions. For IBL and IBL-whisker we report the mean and standard error of the mean of BPS over NN=842 neurons across five test sessions. For Facemap we report the mean and standard deviation of Variance Explained of the first 128 Principal Components of neural activity across five test sessions.
Features IBL (BPS) IBL-whisker (BPS) Facemap (Var Exp) RRR TCN RRR TCN RRR TCN Keypoints $0.169\pm 0.008$ $0.269\pm 0.011$ $0.224\pm 0.047$ $0.403\pm 0.077$ Motion energy $0.086\pm 0.006$ $0.113\pm 0.007$ PCA $0.260\pm 0.010$ $0.309\pm 0.012$ $0.212\pm 0.009$ $0.272\pm 0.011$ $0.177\pm 0.064$ $0.407\pm 0.090$ CEBRA $0.239\pm 0.010$ $0.293\pm 0.012$ $0.204\pm 0.009$ $0.260\pm 0.011$ $0.143\pm 0.046$ $0.278\pm 0.064$ DINOv2 (LVD-142M) $0.183\pm 0.008$ $0.294\pm 0.013$ $0.138\pm 0.007$ $0.269\pm 0.012$ CLIP (WIT-400M) $0.182\pm 0.008$ $0.292\pm 0.012$ $0.131\pm 0.006$ $0.268\pm 0.012$ VideoMAE (Kinetics-400) $0.189\pm 0.010$ $0.318\pm 0.013$ $0.133\pm 0.007$ $0.298\pm 0.013$ ViT-M (IN) $0.192\pm 0.009$ $0.321\pm 0.013$ $0.129\pm 0.007$ $0.301\pm 0.012$ $0.254\pm 0.061$ $0.446\pm 0.100$ ViT-M (IN+PT) $0.172\pm 0.008$ $0.331\pm 0.013$ $0.148\pm 0.007$ $0.311\pm 0.013$ ViT-M (IN+FT) $0.233\pm 0.009$ $0.345\pm 0.014$ $0.211\pm 0.010$ $0.330\pm 0.011$ $0.260\pm 0.051$ $0.461\pm 0.099$ ViT-M (IN+PT+FT) $\mathbf{0.285\pm 0.011}$ $\mathbf{0.347\pm 0.014}$ $0.235\pm 0.010$ $0.328\pm 0.013$ ViT-C (IN+PT) $0.137\pm 0.008$ $0.314\pm 0.013$ $0.120\pm 0.006$ $0.283\pm 0.011$ Beast (IN+PT) $0.205\pm 0.008$ $0.292\pm 0.012$ $0.133\pm 0.006$ $0.309\pm 0.013$ Beast (IN+FT) $0.264\pm 0.010$ $\mathbf{0.347\pm 0.014}$ $\mathbf{0.240\pm 0.010}$ $\mathbf{0.331\pm 0.013}$ $\mathbf{0.281\pm 0.054}$ $\mathbf{0.464\pm 0.089}$ Beast (IN+PT+FT) $0.282\pm 0.011$ $\mathbf{0.347\pm 0.014}$ $0.234\pm 0.010$ $0.326\pm 0.013$